OpenStudy (anonymous):

Simplify, 1.5m^7(-4m^5)^2 24m^17 ya?

OpenStudy (anonymous):

-4^2 = ???

OpenStudy (anonymous):

16

OpenStudy (anonymous):

\[(-4)^2 = -4 \times -4 \]

OpenStudy (anonymous):

80

OpenStudy (anonymous):

waittwait

OpenStudy (anonymous):

oh 12

OpenStudy (anonymous):

12m17

OpenStudy (anonymous):

ya?

OpenStudy (jiteshmeghwal9):